OpenStudy (faman39):

A bag contains 1 blue, 2 white, and 2 red marbles. Match the columns. P( red and red) with replacement P ( white and white ) without replacement P( blue and red) with replacement P( white and red) without replacement 10% 16% 8% 20%

OpenStudy (experimentx):

do you know to find the simple probability?? like probability of getting red marble at first draw??

OpenStudy (unklerhaukus):

P(red)=2/5 P(red,red)=2/5 x 2/5

OpenStudy (unklerhaukus):

P(w,w')=2/5 x 1/4

OpenStudy (unklerhaukus):

or you could just work it out by ordering them in terms of which is most likely